The sum of three numbers is 15. The sum of twice the first​ number,4 times the second​ number, and 5 times the third number is

52. The difference between 6 times the first number and the second number is 40. Find the three numbers.

first​ number:

second​ number:

third​ number:
Answer #1

Noro zutul noty a 2 = 15 - 2n+ 4y + 52252 - Gn-y=40 y = 62-40 - puttry egh M in 6 2 22tylon-yo) +52-52 2n +242-160 + 5z = 52In case of any doubt please ask in comment section.